We investigate the integer sequence (t_{n})_{n∈ℤ} defined by t_{n} = 0 if n ≤ 0, t_{1} = 1, and t_{n} = ∑_{i=1}^{n−1} t_{n−t_{i}} for n ≥ 2. This sequence has the following properties: if we consider f_{n}(X) := −1+∑_{i=1}^{n} X^{t_{i}} and take x_{n} to be the real positive number such that f_{n}(x_{n}) = 0, then

lim_{n→∞} \frac{t_{n}}{t_{n+1}} = lim_{n→∞} x_{n} = 0.410098516 ⋯

Moreover, if u is the real positive number such that 1 = ∑_{i=1}^{∞} u^{−t_{i}}, then there is a positive constant M such that t_{n} ∼ Mu^{n}.